OpenStudy (anonymous):

NEED ANSWER FAST WILL MEDAL Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

lol why needed fast? XD

zepdrix (zepdrix):

Our point-slope form of a line looks like this: \(\rm y-y_1=m(x-x_1)\) Oh, let's not worry about that just yet though, I guess finding slope is more important here. \[\large\rm m=\frac{y_2-y_1}{x_2-x_1}\]And you were given two points,\[\large\rm (x_1,y_1)=(-3,2),\qquad\qquad\qquad (x_2,y_2)=(2,1)\]

zepdrix (zepdrix):

Understand how to find the slope value? Plug in the stuff.

OpenStudy (anonymous):

Becauase i need to finish math today so I dont fail, I started late so it was harder to catch up

zepdrix (zepdrix):

Grr OpenStudy keeps crashing on me >;c

OpenStudy (mathmale):

I'm sorry you're under pressure. Your best bet is to master the basics; "slope of a straight line" is one of those basics. See zepdrix' previous post; he gave you the correct formula for "slope of a straight line." Any questions about this? If not, please take the given data (the 2 points) and use this slope formula to find the slope of the line connecting these two points.

OpenStudy (mathmale):

Important: Please show your work.

OpenStudy (anonymous):

Is it c